A fish is little below the surface of water with critical angle 49degree find the angular vision of fish

Grade:12

1 Answers

Arun
6 years ago
Dear student
If we draw the figure,
We get
\theta = 2c
 = 2 *49° = 98°
Regards
Arun (askIITians forum expert)